# Looking Right is Sometimes Right: Investigating the Capabilities of Decoder-only LLMs for Sequence Labeling

### v2

@@ Abstract
Pre-trained language models based on masked language modeling (MLM) excel in natural language understanding (NLU) tasks. While fine-tuned MLM-based encoders consistently outperform causal language modeling decoders of comparable size, recent decoder-only large language models (LLMs) perform on par with smaller MLM-based encoders. Although their performance improves with scale, LLMs fall short of achieving state-of-the-art results in information extraction (IE) tasks, many of which are formulated as sequence labeling (SL). We hypothesize that LLMs’ poor SL performance stems from causal masking, which prevents the model from attending to tokens on the right of the current token. Yet, how exactly and to what extent LLMs’ performance on SL can be improved remains unclear. We explore techniques for improving the SL performance of open LLMs on IE tasks by applying layer-wise removal of the causal mask (CM) during LLM fine-tuning. This approach yields performance gains competitive with state-of-the-art SL models, matching or outperforming the results of CM removal from all blocks. Our findings hold for diverse SL tasks, demonstrating that open LLMs with layer-dependent CM removal outperform strong MLM-based encoders and even instruction-tuned LLMs.
@@ 1 Introduction
In spite of LLMs’ good performance on NLU tasks, there is still considerable room for improvement, even for the largest decoder models, such as ChatGPT, which fall far behind SOTA results on fundamental NLP tasks. This holds particularly for information extraction (IE) tasks (Han et al., 2023), such as named entity recognition (NER), aspect-based sentiment analysis (ABSA), and event extraction (EE). These tasks are often formulated as sequence labeling (SL) tasks, and tackling SL by prompting LLMs proved quite difficult (Wang et al., 2023a). However, it is unclear how exactly and to what extent the subpar performance of LLMs on SL tasks can be remedied. As the community stopped scaling up encoders, LLMs are solidifying their position as field’s de facto standard. Thus, improving LLMs’ SL performance is the sole viable option.
@@
The poor performance of LLMs on SL tasks can be traced back to the information encoded in the hidden states. Using the hidden state of the decoder’s last token serves as a reliable feature for classifying the input sequence. However, the same tactic does not suffice for SL because the causal mask (CM), integral to endowing decoders with text generation capabilities, limits bidirectional information flow, preventing the model from “looking right,” i.e., attending to tokens in the sequence positioned to the right of the current token. For many SL tasks, the token’s label depends on the succeeding tokens, and omitting this context often results in subpar performance. Recently, Li et al. (2023) found that completely removing the CM from the Llama2 model during fine-tuning substantially enhances SL performance on NER.
In this paper, we explore techniques to enhance the SL performance of open LLMs on IE tasks. Building on insight from Li et al. (2023), we hypothesize that removing the CM from all decoder layers may not be beneficial for all SL tasks. To confirm this, we experiment with layer-wise CM removal across LLM blocks, observing gains competitive with SOTA models, in contrast to removing or keeping the CM in all layers. We extend our analysis to a series of IE tasks (NER, ABSA, and EE), and demonstrate that layer-wise “looking right” is highly task-dependent. We compare against strong encoder-only sequence taggers and instruction-tuned LLMs. Open LLMs without CMs in specific layers outperform these baselines in almost all scenarios. Finally, we compare with encoder-only models, assessing the relative strengths of the decoder and encoder architectures regarding parameter scale, training data, and the particular SL task. We pre-train small encoder and decoder models from scratch with an identical number of parameters on the same data. Our findings indicate the superiority of the same-scale encoder architecture over the decoder on SL tasks, which is consistent even when the CM is removed while fine-tuning the small pre-trained decoders. This suggests that the observed SL performance gain in LLMs upon CM removal stems from scaling up.
@@ 2 Background and Related Work
Pre-training with CLM teaches the model to generate text and does not allow it to use the sequence context right to the token being processed. This is enabled via a CM, which prevents the model from attending to future tokens. While beneficial and necessary for coherent text generation, restricting the model to look right is detrimental if one wants to use the decoder to produce fully contextualized token-level embeddings with the decoder. Although decoder-only models perform poorly on sequence labeling tasks (i.e., tasks where each individual token is assigned a label), they perform quite well on sequence classification tasks (i.e., tasks where the entire sequence is assigned one label). Therefore, many decoder models implement sequence classification heads on top of decoder blocks, but not SL heads (Wolf et al., 2020).
@@
Many IE tasks, such as NER, subtasks of ABSA, and EE, can effectively be formulated as SL tasks. Traditionally, these tasks were addressed with encoders, mapping each token’s hidden state to task label space with a linear layer on top of the last encoder block (Pontiki et al., 2015; Zhang et al., 2020; Zhou and Chen, 2021). However, this approach necessitates training a separate model for each task. More recent methods aim to infer the underlying structure of multiple IE tasks simultaneously and develop universal information extraction (UIE) models, employing encoder, decoder, or encoder-decoder architectures (Paolini et al., 2021; Lu et al., 2022; Fei et al., 2022; Wang et al., 2022a; Ping et al., 2023; Zhu et al., 2023; Lou et al., 2023). While the practical advantages of UIE are obvious, including more IE tasks would still require re-training the UIE extractors. Therefore, we stick to the one model per SL task strategy, relying on parameter-efficient methods that yield compact and extensible models (Houlsby et al., 2019).
@@
ICL is a powerful method for scrutinizing LLMs, employed by prompting LLMs with task demonstrations. IE tasks can be tackled with ICL, relying on the emergent properties of LLMs. Nevertheless, Pang et al. (2023) report that the performance of LLMs leveraging ICL lags behind the SOTA results of supervised IE models. They develop a model that learns to guide the prompt for improved LLM performance. Another notable example is GPT-NER (Wang et al., 2023a), transforming the NER task into a generation one utilizing ICL and ChatGPT.
@@ 4.1 Sequence Labeling Tasks
Table 1: Statistics for the datasets and their splits. We show number of sentences per split, total number of sentences, and right-side dependency relations ratio.
@@ 4.2 Training Details and Hyperparameters
We use Llama2-7B and Mistral-7B with CLM heads for IT. Training is done with QLoRA based on cross-entropy loss. All hyperparameters are inherited from SL experiments except for the batch size. The batch size depends on the maximum sequence length for packing the dataset examples. Short examples are packed together in the same input to increase training efficiency. This length is set to 512 for the Rest14 dataset and 1024 for other datasets since Rest14 has shorter instruction examples on average and has the smallest training set (cf. Table 1). We observe that ATE+ATP models need double the epochs (10 epochs) to learn the task. We hypothesize this is due to the low number of training examples compared to other datasets.
@@ 5.3 Comparison with SOTA IE Models
Our results are competitive with SOTA. However, a fair comparison is challenging due to differences in training and evaluation. For example, the SOTA F1 score for CoNLL03 NER, reported by Wang et al. (2021), is $0.946$. This result, however, was obtained using a model trained on the merged training and validation sets. Moreover, important details of the evaluation setup were not specified (micro vs. macro F1 score, token- vs. span-based evaluation, and evaluating the model prediction on the first token of tokenized words from the input sequence vs. on all tokens). Further, the authors whose work we build upon, Li et al. (2023), report SOTA results on CoNLL03 NER. The score they report is $0.932$, which is competitive, but not SOTA. Upon code inspection, we found that they truncate sequences longer than 64 tokens to fit the data into GPU memory. In contrast, we managed to keep the sequence length at 128 tokens. These decisions have a significant impact on overall performance. Scaria et al. (2023) achieve SOTA F1 score of $0.928$ on Rest14 dataset and ATE task with IT. Yang and Li (2021) report SOTA macro F1 of $0.863$ on ATP task with DeBERTa model He et al. (2020). The SOTA F1 score of $69.8$ on ACE05 for TC is achieved by Wang et al. (2022a).
@@ 6 Conclusion
The landscape of PLMs shifted from encoder to decoder dominance in various NLU tasks. While large decoder models show remarkable performance even without fine-tuning, challenges persist in IE tasks, typically formulated as sequence labeling. In a series of experiments, we showed that LLMs can yield performance competitive with SOTA on sequence labeling IE tasks when the causal mask is removed from specific decoder blocks. Future work should should investigate CM removal for individual layers, analyze task-specific differences in CM removal, and investigate the options for predicting the impact of CM removal without requiring extensive fine-tuning. Comparisons with encoders stress the significance of architecture and scale in SL tasks. A noticeable gap exists between small-scale encoder and decoder models, where decoder models do not benefit from CM removal at a small scale.
@@ 7 Limitations
In our experiments, we ensure reliability by averaging performance scores over five runs with different random seeds. Increasing the sample size for averaging would enhance reliability further. However, maintaining fixed learning rates and other hyperparameters across experiments might have led to suboptimal adaptation for sequence labeling tasks. Exploring additional unmasking configurations could provide valuable insights and potential improvements. While numerous open LLMs are available, we focus solely on Llama2 and Mistral, both with 7B parameters. Enhancing prompt templates used for instruction tuning could boost overall instruction tuning performance. Given more computing resources, experimenting with larger models would be feasible. Additionally, our experiments were limited to English-language datasets. Extending the analysis to sequence labeling tasks in other languages and incorporating more diverse datasets could yield further insights.

### v3

@@ Abstract
Pre-trained language models based on masked language modeling (MLM) excel in natural language understanding (NLU) tasks. While fine-tuned MLM-based encoders consistently outperform causal language modeling decoders of comparable size, recent decoder-only large language models (LLMs) perform on par with smaller MLM-based encoders. Although their performance improves with scale, LLMs fall short of achieving state-of-the-art results in information extraction (IE) tasks, many of which are formulated as sequence labeling (SL). We hypothesize that LLMs’ poor SL performance stems from causal masking, which prevents the model from attending to tokens on the right of the current token. Yet, how exactly and to what extent LLMs’ performance on SL can be improved remains unclear. We explore techniques for improving the SL performance of open LLMs on IE tasks by applying layer-wise removal of the causal mask (CM) during LLM fine-tuning. This approach yields performance gains competitive with state-of-the-art SL models, matching or outperforming the results of CM removal from all blocks. Our findings hold for diverse SL tasks, demonstrating that open LLMs with layer-dependent CM removal outperform strong MLM-based encoders and even instruction-tuned LLMs. [^1] Code: https://github.com/dd1497/llm-unmasking.
@@
$\dagger$\$$\dagger$\$footnotetext: Corresponding author: david.dukic@fer.hr
@@ 1 Introduction
Figure 1: Layer-wise causal mask removal from decoder block groups in a decoder-only LLM. Here, the causal mask is removed from the top eight decoder blocks of the Llama2-7B model to enable bidirectionality during fine-tuning, which proves beneficial for many SL tasks.
Despite LLMs’ good performance on NLU tasks, there is still considerable room for improvement, even for the largest decoder models, such as ChatGPT, which fall far behind SOTA results on fundamental NLP tasks. This holds particularly for information extraction (IE) tasks (Han et al., 2023), such as named entity recognition (NER), aspect-based sentiment analysis (ABSA), and event extraction (EE). These tasks are often formulated as sequence labeling (SL), and tackling SL by prompting LLMs proved quite difficult (Wang et al., 2023a). However, it is unclear how exactly and to what extent the subpar performance of LLMs on SL tasks can be remedied. [^3] Although LLMs can solve many NLU tasks, including, but not limited to, IE tasks used for the knowledge base population, they have their limitations. We still need specialized methods for SL-based IE where we want to identify and count information elements in the text. This is especially useful for computational social science applications. As the community stopped scaling up encoders, LLMs are solidifying their position as the field’s de facto standard. Thus, improving LLMs’ SL performance is the sole viable option.
@@
The poor performance of LLMs on SL tasks can be traced back to the information encoded in the hidden states. Using the hidden state of the decoder’s last token serves as a reliable feature for classifying the input sequence. However, the same tactic does not suffice for SL because the causal mask (CM), integral to endowing decoders with text generation capabilities, limits bidirectional information flow, preventing the model from “looking right,” i.e., attending to tokens in the sequence positioned to the right of the current token. For many SL tasks, the token’s label depends on the succeeding tokens, and omitting this context often results in subpar performance. Recently, Li et al. (2023) found that completely removing the CM from the Llama2 model during fine-tuning substantially enhances SL performance on NER. Others also experimented with the CM removal and enabling bidirectionality concurrently with our work (BehnamGhader et al., 2024; Lee et al., 2024).
In this paper, we explore techniques to enhance the SL performance of open LLMs on IE tasks. Building on insight from Li et al. (2023), we hypothesize that removing the CM from all decoder layers may not be beneficial for all SL tasks. To confirm this, we experiment with layer-wise CM removal across LLM blocks, observing gains competitive with SOTA models, in contrast to removing or keeping the CM in all layers. We extend our analysis to a series of IE tasks (NER, ABSA, and EE) and demonstrate that layer-wise “looking right” is highly task-dependent. We compare against strong encoder-only sequence taggers and instruction-tuned LLMs. Open LLMs without CMs in specific layers outperform these baselines in almost all scenarios. Finally, we compare with encoder-only models, assessing the relative strengths of the decoder and encoder architectures regarding parameter scale, training data, and the particular SL task. We pre-train small encoder and decoder models from scratch with an identical number of parameters on the same data. Our findings indicate the superiority of the same-scale encoder architecture over the decoder on SL tasks, which is consistent even when the CM is removed while fine-tuning the small pre-trained decoders. This suggests that the observed SL performance gain in LLMs upon CM removal stems from scaling up.
@@ 2 Background and Related Work
Pre-training with CLM teaches the model to generate text and does not allow it to use the sequence context right to the token being processed. This is enabled via a CM, which prevents the model from attending to future tokens. While beneficial and necessary for coherent text generation, restricting the model to look right is detrimental if one wants to use the decoder to produce fully contextualized token-level embeddings with the decoder. Although decoder-only models perform poorly on sequence labeling tasks (i.e., tasks where each individual token is assigned a label), they perform quite well on sequence classification tasks (i.e., tasks where the entire sequence is assigned one label). Therefore, many decoder models implement sequence classification heads on top of decoder blocks but not SL heads (Wolf et al., 2020).
@@
Many IE tasks, such as NER, subtasks of ABSA, and EE, can effectively be formulated as SL tasks. Traditionally, these tasks were addressed with encoders, mapping each token’s hidden state to task label space with a linear layer on top of the last encoder block (Pontiki et al., 2015; Zhang et al., 2020; Zhou and Chen, 2021). However, this approach necessitates training a separate model for each task. More recent methods aim to infer the underlying structure of multiple IE tasks simultaneously and develop universal information extraction (UIE) models, employing encoder, decoder, or encoder-decoder architectures (Paolini et al., 2021; Lu et al., 2022; Fei et al., 2022; Wang et al., 2022a; Ping et al., 2023; Zhu et al., 2023; Lou et al., 2023; Ding et al., 2024). While the practical advantages of UIE are obvious, including more IE tasks would still require re-training the UIE extractors. Therefore, we stick to the one model per SL task strategy, relying on parameter-efficient methods that yield compact and extensible models (Houlsby et al., 2019).
@@
ICL is a powerful method for scrutinizing LLMs, employed by prompting LLMs with task demonstrations. IE tasks can be tackled with ICL, relying on the emergent properties of LLMs. Nevertheless, Pang et al. (2023) report that the performance of LLMs leveraging ICL lags behind the SOTA results of supervised IE models. They develop a model that learns to guide the prompt for improved LLM performance. Another notable example is GPT-NER (Wang et al., 2023a), transforming the NER task into a generation one utilizing ICL and ChatGPT, while Blevins et al. (2023) and Mehta et al. (2024) apply prompt-based methods to structured prediction tasks such as NER and semantic role labeling.
@@ 4.1 Sequence Labeling Tasks
Table 1: Statistics for the datasets and their splits. We show the number of sentences per split, the total number of sentences, and the right-side dependency relations ratio (RDRR).
@@ 4.2 Training Details and Hyperparameters
We use Llama2-7B and Mistral-7B with CLM heads for IT. Training is done with QLoRA based on cross-entropy loss. All hyperparameters are inherited from SL experiments except for the batch size. The batch size depends on the maximum sequence length for packing the dataset examples. Short examples are packed together in the same input to increase training efficiency. This length is set to 512 for the Rest14 dataset and 1024 for other datasets since Rest14 has shorter instruction examples on average and has the smallest training set (cf. Table 1). We observe that ATE+ATP models need double the epochs (10 epochs) to learn to extract aspect terms and their polarity, and we hypothesize that this is due to the low number of training examples compared to other datasets (~3k vs. ~14k).
@@ 5.3 Comparison with SOTA IE Models
Our results are competitive with SOTA. However, a fair comparison is challenging due to differences in training and evaluation. For example, the SOTA F1 score for CoNLL03 NER, reported by Wang et al. (2021), is $0.946$. This result, however, was obtained using a model trained on the merged training and validation sets. Strong results on CoNLL03 NER were reported by Liu et al. (2022), reaching an F1 score of $0.941$ without task-specific feature engineering, relying solely on a conditional language model with explicit modeling of the target structure. Further, the authors whose work we build upon, Li et al. (2023), report SOTA results on CoNLL03 NER. The score they report is $0.932$, which is competitive, but not SOTA. Upon code inspection, we found that they truncate sequences longer than 64 tokens to fit the data into GPU memory. In contrast, we managed to keep the sequence length at 128 tokens. These decisions have a significant impact on overall performance. Although the reported results in related work are SOTA or close to SOTA, these studies are representative of a common problem in the field, namely the fact that the important evaluation details are not always communicated properly (micro vs. macro F1 score, token- vs. span-based evaluation, and evaluating the model prediction on the first token of tokenized words from the input sequence vs. on all tokens). Scaria et al. (2023) achieve SOTA F1 score of $0.928$ on Rest14 dataset and ATE task with IT. Yang and Li (2021) report SOTA macro F1 of $0.863$ on ATP task with DeBERTa model He et al. (2020). The SOTA F1 score of $69.8$ on ACE05 for TC is achieved by Wang et al. (2022a).
@@ 6 Conclusion
The landscape of PLMs shifted from encoder to decoder dominance in various NLU tasks. While large decoder models show remarkable performance even without fine-tuning, challenges persist in IE tasks, typically formulated as sequence labeling. In a series of experiments, we showed that LLMs can yield performance competitive with SOTA on sequence labeling IE tasks when the causal mask is removed from specific decoder blocks. Future work should investigate CM removal for individual layers, analyze task-specific differences in CM removal, and investigate the options for predicting the impact of CM removal without requiring extensive fine-tuning. Promising alternatives to trying out all the unmasking configurations could involve leveraging dynamic programming or estimating a network’s final performance without retraining, for example, using methods that rely on the Fisher information, similar to the FIT method by Zandonati et al. (2022). Additionally, performance estimation for CM removal from a layer could be improved using representation similarity analysis methods such as centered kernel alignment Kornblith et al. (2019). These approaches might reduce the number of unmasking configurations that need to be evaluated to find the best one. Comparisons with encoders stress the significance of architecture and scale in SL tasks. A noticeable gap exists between small-scale encoder and decoder models, where decoder models do not benefit from CM removal at a small scale.
@@ 7 Limitations
In our experiments, we ensure reliability by averaging performance scores over five runs with different random seeds. Increasing the sample size for averaging would enhance reliability further. However, maintaining fixed learning rates and other hyperparameters across experiments might have led to suboptimal adaptation for sequence labeling tasks. Exploring additional unmasking configurations could provide valuable insights and potential improvements. While numerous open LLMs are available, we focus solely on Llama2 and Mistral, both with 7B parameters. Enhancing prompt templates used for instruction tuning could boost overall instruction tuning performance. Given more computing resources, experimenting with larger models would be feasible. Additionally, our experiments were limited to English-language datasets. Extending the analysis to sequence labeling tasks in other languages and incorporating more diverse datasets could yield further insights. Finally, the main limitation of our work is the fact that one needs to try out all the possible unmasking configurations to find the best one on the validation set. Trying out all the possible configurations requires extensive fine-tuning.